### v1

@@ Abstract
High 1/f noise in CryoCMOS devices is a critical parameter to keep under control in the design of complex circuits for low temperatures applications. Current models predict the 1/f noise to scale linearly with temperature, and gate oxide defects are expected to freeze out at cryogenic temperatures. Nevertheless, it has been repeatedly observed that 1/f noise deviates from the predicted behaviour and that gate oxide defects are still active around 4.2 K, producing random telegraph noise. In this paper, we probe single gate oxide defects in 2500 nMOS devices down to 5 K in order to investigate the origin of 1/f noise in CryoCMOS devices. From our results, it is clear that the number of defects active at cryogenic temperatures resulting in random telegraph noise is larger than at 300 K. Threshold voltage shifts due to charged defects are shown to be exponentially distributed, with different modalities across temperatures and biases: from monomodal at 300 K to trimodal below 100 K. The third mode is interpreted in the framework of percolation theory. By fitting these distributions, it is shown that more than 80% of the detected defects belongs to the oxide bulk. Afterwards, starting from the raw data in time domain, we reconstruct the low frequency noise spectra, highlighting the contributions of defects belonging to different branches and, therefore, to different oxide layers. This analysis shows that, although interface traps and large defects associated with the third mode are the main sources of 1/f noise at 5 K, bulk oxide defects still contribute significantly to low-frequency noise at cryogenic temperatures. Finally, we show that defect time constants and step heights are uncorrelated, proving that elastic tunnelling model for charge trapping is not accurate.
@@
In recent years, technologies related to quantum computing progressed considerably, in terms of both qubits quality and performance [1] and of co-integration with classical electronics for qubits control and read-out [2, 3]. Since in most of the present technologies the control read-out electronics is placed at room temperature, outside the cryogenically cooled environment (dilution refrigerator) of the qubits, and each qubit must be addressed individually through long wires, the number of qubits that can be integrated on a single chip and tested is severely limited. Bringing control circuitry into the cryogenic environment has become a much-awaited solution - yet this comes with its own challenges and limitations. One of the main challenges in the design of complex cryogenic circuits is the lack of process design kits (PDKs) at low temperatures which, in turns, mirrors a lack of understanding and modelling of MOSFETs at liquid He temperature and below. Although in the last decade several works enlightened many critical aspects of cryogenic devices’ static operation [4, 5, 6, 7], the understanding of defects dynamics and the resulting reliability issues requires more effort [8, 9]. In particular, low frequency noise (LFN) was shown to not scale linearly with temperature [10, 11, 12, 13], as predicted by a variety of models that refer to carrier number fluctuations (CNFs) with correlated mobility fluctuations (MFs) [14]. This "excess 1/f noise" was proposed to be caused by density of states (DOS) bandtails [10], which are shown to be responsible for the subthreshold slope saturation at cryogenic temperatures [5, 7, 15]. However, understanding the origin of this excess noise is extremely challenging, since LFN is generally measured on large area devices and therefore it provides the average response of thousands of defects, making the characterization of individual traps impractical. For this reason, single defect spectroscopy techniques, such as random telegraph noise (RTN), were recently employed on small area devices to get insights on trap physics at cryogenic temperatures [9, 16, 17]. In some studies, RTN was attributed to localized band-edge states [16] or to interface defects [18], whereas in [17] it was shown that RTN can also originate from defects far from the interface. In most RTN works, authors focused on the characterization of few selected devices, and the conclusions obtained can be hardly generalized to a statistically significant ensemble of defects. In other words, it is not possible to fully comprehend the statistical physical behaviour of a specific technology nor to link the activity of few defects with the unexpected 1/f behaviour of large area devices. To bridge this gap, we therefore use large-scale statistics of individual devices. We employ random telegraph noise characterization on an entire array of 2500 n-type MOSFETs. Using the ensemble statistics, we obtain distributions of single trap properties that can be linked to the physical mechanisms responsible for the 1/f noise across temperatures.
@@
This work is structured as follows. First, RTN amplitudes are extracted using a state-of-the-art Machine Learning (ML) algorithm together with a Hidden Markov Model (HMM) algorithm. Afterwards, their distributions across temperatures and overdrives are plotted and extensively discussed. Threshold voltage shifts due to charged defects are exponentially distributed according to different modalities at different temperatures. By fitting these distributions, the average threshold voltage shift together with the statistical defect locations are extracted. Original time traces are then used to reconstruct the low frequency noise spectra of the entire array, which follows a 1/f behaviour down to cryogenic temperatures. The LFN is higher at K, in excellent agreement with the observations already reported in literature [10, 11, 13]. The contribution to 1/f noise of defects coming from different modes is further explored. Finally, the strong correlation between time constants and RTN amplitudes predicted by the elastic tunnelling model is tested [19].
@@ 2.1 Chip design and characterization
An array of 2500 nominally identical and minimally-spaced nanoscale nMOS devices (WxL = 90 nm x 28 nm) was used in this work (Fig. 1(a)). It consists of 2500 individually-addressable gate lines, each connected to 10 devices. The structure was designed and fabricated using a commercial 28 nm high-k metal gate CMOS technology (see Methods). The chip was bonded on a custom designed PCB and placed in a Lakeshore CPX-LVT cryogenic probe station, equipped with a custom designed 48-lines break-out box [20]. The measurements were performed on specifically designed hardware, which allowed to not only program and measure the array but also to change the sample temperature, enabling semi-automated measurement sequences.
The array was tested at six different temperatures, from T=300 K down to T=5 K. At each temperature, the characteristics in linear regime ( V) of the entire device ensemble were collected (see Fig. 1(b)). This step was used to remove the broken devices from the analysis, extracting the threshold voltage of every working device and map the data to . The threshold voltage was calculated by using the Lambert-W function method [21], which was shown to be effective down to liquid helium temperature [22]. Afterwards, the drain current of each device was recorded for a total recording time of , with a sampling time of ms, at five gate bias points: from to V, with a step V. The drain bias was fixed at V.
In Fig. 1 two examples of traces showing RTN are reported. In the first case (Fig. 1(c)), the RTN is produced by a single two-levels defect, which captures (low ) and emits (high ) individual carriers from the channel. The second case, instead, represents a more complicated signal, with two two-level defects active at the same time. A key aspect of this work is the development and the employment of an algorithm able to treat multiple defect signals. Indeed, this need turned out to be crucial for the RTN analysis at cryogenic temperatures, where the number of traces containing two defects was around 30% of the total.
To extract the defect average step height , as well as the average emission and capture time constants , a method based on a physics-informed machine learning (ML) algorithm [23] coupled with a hidden Markov model (HMM) [24] was developed. The ML algorithm allowed to automatically detect RTN in the traces, distinguishing between one or multi-defect signals. In case of single-defect RTN, and were directly extracted from the signal reconstructed by the ML algorithm, whereas, in case of two active defects, the raw data were further analysed using the HMM. The benefits of this method are manifold: it requires little manual interaction, so it is suited for the treatment of large amount of data; given its complexity, it is reasonably fast (it can run on a personal computer); it allows the analysis of multi-defects RTN, as previously discussed. Further details are provided in the Methods section.
@@ 2.2 Defects’ step heights distributions
In Fig. 2(a) are plotted the CCDFs of the step heights extracted from the measurement data of approximately 2500 devices at both K and K, and V. With decreasing temperature, the shape of the CCDF changes. To determine the modality of each exponential distribution in a statistically rigorous manner, a fitting method based on the maximization of the likelihood function was developed together with an outliers detection algorithm (see Supplement Material). Furthermore, this fitting procedure allows the extraction of and (see eq.(1)). It is worth pointing out that includes also the undetected defects (more details can be found in the Supplement Material). At K the CCDF is monomodal, although the technology used in this work consists of devices with a stack. Indeed, in case of multi-layers stacks, single-defects step heights generally follow a bimodal distribution: the low- mode belongs to bulk defects, whereas the high- mode is attributed to interface defects [31, 35]. This division is based on electrostatic considerations: defects are smaller compared to ones, since they are more far away from the channel and their electrostatic impact is lower. In the present case, most of defects probably have a smaller than the detection limit of both the ML algorithm and the HMM used in the time traces analysis (see Supplement Materials). At K and V, the detection limit was set to mV. It is worth pointing out that this argument finds confirmation in Fig. 2(d), where the extracted values of are plotted. Even at K, where is generally found to be higher, as it will be discussed later, mV. Therefore, the CCDF shown in Fig. 2(a) belongs only to defects. At this point, it can be instructive to compare the experimental with predicted by the CSA. Coherently with the literature [28, 29, 33], mV is more than two times larger than mV. As the temperature decreases to K, the CCDF still appears to be monomodal (see Supplement Materials), and room temperature considerations hold.
@@
Starting from K (see Supplement Material), and down to K (Fig. 2(a)), distributions change modality, with the appearance of two other branches: one at small and another one at very large . The three modes are highlighted in Fig. 2(b), where the probability density function (PDF) of the trimodal exponential fit is plotted. This latter is given by the sum of three independent monomodal exponential PDFs, each one being described by a different set of and . The first mode is attributed to defects, which are now detectable because of both the reduced background noise and the larger transistor’s . Indeed, since , small threshold voltage shifts are easier to be detected at cryogenic temperatures because they result in a larger . The second mode is linked to defects, which enables a direct investigation of their temperature scaling from K up to K, as discussed later. Finally, a third mode is observed, which manifests itself predominantly for large .
@@
A defect with a given amplitude is associated with each mode with a certain probability. In Fig. 2(b), there are two values that have the same probability to belong to two modes: these are given by the intersection points and . A threshold voltage shift due to charged defects lying on the left side of , for example, is more likely to fall in the branch. The probability that the amplitude of a detected oxide defect falls into intervals of values defined by and is calculated using the cumulative distribution function (CDF) and it is plotted in Fig. 2(c). More than 80% of measured defects have mV, and more than 60% of these belong to . Therefore, contrary to what has been previously published on RTN at cryogenic temperatures [13] [16], most of the activity seems to come from bulk defects, which do not freeze out. In Fig. 2(d) the temperature dependence of at V is plotted. Globally, increases at cryogenic temperatures, independently on the branch considered. Defects linked to the third mode, however, appear to be more sensitive to temperature scaling. and defects also show an increase in their average threshold voltage shift as the temperature is lowered. As it will be discussed below, this trend of can be effectively interpreted in the framework of percolation theory.
@@
In order to interpret the shape of the CCDF at cryogenic temperatures (see Fig. 2(a) and Fig. 2(b)), it is worth mentioning that, in the field of cryogenic electronics, percolation theory was proposed in the late ’70s to interpret the smooth change from metallic-like to thermally-activated transport in MOSFET’s inversion layer when it operates around and below and at K [36, 37, 38]. This metal-insulator transition [39] is generally attributed to surface potential fluctuations caused by interface disorder. Macroscopically, these fluctuations are usually modelled as an additional density of states below the conduction (valence) band edge [5, 7, 15]. Therefore, percolation theory has been employed independently to describe two apparently different phenomena, both originating from surface potential fluctuations: electrical conduction in band tail states at cryogenic temperatures and statistical distribution of threshold voltage shifts due to charged defects at room temperature. Starting from these considerations, percolation theory is here proposed to interpret the distributions below K (see Fig. 2(a) and Supplement Materials). In this framework, the observed CCDFs are a manifestation of both the increased sensitivity to disorder and the appearance of interface localized states. Fluctuations in surface potential mirrors in a non uniform distribution of inversion carriers along the channel. If they have enough thermal energy, the main consequence of such fluctuations is a non homogeneous current flow from source to drain which, as previously mentioned, is responsible for the exponential distribution of s. As temperature drops, carriers thermal energy decreases accordingly, and potential fluctuations can give rise to regions where carriers are energetically prevented to flow. It is crucial to highlight that, at room temperature, the channel is modulated by disorder but it is continuous. At cryogenic temperatures, instead, the channel is a mixture of conductive paths and insulating regions [36]. This increased channel inhomogeneity amplifies the electrostatic impact of all defects at low temperatures, in agreement with the trend previously discussed and plotted in Fig. 2(d). RTN fluctuations producing the third branch in Fig. 2(a) are therefore interpreted as the result of the activity of those defects whose electrostatic impact is particularly enhanced by the larger channel inhomogeneity. On the other hand, especially around K, potential fluctuations can give rise to localized states, which contribute to transport through thermally-activated and also tunnelling processes. Some of the large in Fig. 2(a) could be caused by the activity of these localized states as well, but further analyses are required to confirm that.
@@
Increasing , the modality of the CCDF changes, passing from trimodal to monomodal, as shown in Fig. 3(a). In this case, the third branch in the CCDF starts to vanish, since the channel gets stronger and percolation transport is mitigated. Furthermore, also the small mode disappeared. This can be attributed to both a reduction in measurement sensitivity and to a shift in the active energy area scanned. Concerning the first argument, at higher it is harder to detect small defects, since is larger. Moreover, changing the gate bias, the Fermi level changes as well, and the energy area of the defect bands swept changes accordingly. Therefore, at V, CCDFs are attributed only to the activity of defects in the layer (see also Supplement Material).
Finally, in Fig. 3(b), is plotted as function of .

### v2

@@ Abstract
High 1/f noise in CryoCMOS devices is a critical parameter to keep under control in the design of complex circuits for low temperatures applications. Current models predict the 1/f noise to scale linearly with temperature, and gate oxide defects are expected to freeze out at cryogenic temperatures. Nevertheless, it has been repeatedly observed that 1/f noise deviates from the predicted behaviour and that gate oxide defects are still active around 4.2 K, producing random telegraph noise. In this paper, we probe single gate oxide defects in 2500 nMOS devices down to 5 K in order to investigate the origin of 1/f noise in CryoCMOS devices. From our results, it is clear that the number of defects active at cryogenic temperatures resulting in random telegraph noise is larger than at 300 K. Threshold voltage shifts due to charged defects are shown to be exponentially distributed, with different modalities across temperatures and biases: from monomodal at 300 K to trimodal below 100 K. Distributions’ shapes are interpreted in the framework of percolation theory. By fitting these distributions, it is shown that more than 80% of the detected defects belongs to the oxide bulk. Afterwards, starting from the raw data in time domain, we reconstruct the low frequency noise spectra, highlighting the contributions of defects belonging to different branches and, therefore, to different oxide layers. This analysis shows that, although interface traps and large defects associated with the third mode are the main sources of 1/f noise at 5 K, bulk oxide defects still contribute significantly to low-frequency noise at cryogenic temperatures. Finally, we show that defect time constants and step heights are uncorrelated, proving that elastic tunnelling model for charge trapping is not accurate.
@@
In recent years, technologies related to quantum computing progressed considerably, in terms of both qubits quality and performance [1] and of co-integration with classical electronics for qubits control and read-out [2, 3]. Since in most of the present technologies the control read-out electronics is placed at room temperature, outside the cryogenically cooled environment (dilution refrigerator) of the qubits, and each qubit must be addressed individually through long wires, the number of qubits that can be integrated on a single chip and tested is severely limited. Bringing control circuitry into the cryogenic environment has become a much-awaited solution - yet this comes with its own challenges and limitations. One of the main challenges in the design of complex cryogenic circuits is the lack of process design kits (PDKs) at low temperatures which, in turns, mirrors a lack of understanding and modelling of MOSFETs at liquid He temperature and below. Although in the last decade several works enlightened many critical aspects of cryogenic devices’ static operation [4, 5, 6, 7], the understanding of defects dynamics and the resulting reliability issues requires more effort [8, 9]. In particular, low frequency noise (LFN) was shown to not scale linearly with temperature [10, 11, 12, 13], as predicted by a variety of models that refer to carrier number fluctuations (CNFs) with correlated mobility fluctuations (MFs) [14, 15]. This "excess 1/f noise" was proposed to be caused by density of states (DOS) bandtails [10], which are shown to be responsible for the subthreshold slope saturation at cryogenic temperatures [5, 7, 16]. However, understanding the origin of this excess noise is extremely challenging, since LFN is generally measured on large area devices and therefore it provides the average response of thousands of defects, making the characterization of individual traps impractical. For this reason, single defect spectroscopy techniques, such as random telegraph noise (RTN), were recently employed on small area devices to get insights on trap physics at cryogenic temperatures [9, 17, 18]. In some studies, RTN was attributed to localized band-edge states [17] or to interface defects [19], whereas in [18] it was shown that RTN can also originate from defects far from the interface. In most RTN works, authors focused on the characterization of few selected devices, and the conclusions obtained can be hardly generalized to a statistically significant ensemble of defects. In other words, it is not possible to fully comprehend the statistical physical behaviour of a specific technology nor to link the activity of few defects with the unexpected 1/f behaviour of large area devices. To bridge this gap, we therefore use large-scale statistics of individual devices. We employ random telegraph noise characterization on an entire array of 2500 n-type MOSFETs. Using the ensemble statistics, we obtain distributions of single trap properties that can be linked to the physical mechanisms responsible for the 1/f noise across temperatures.
@@
This work is structured as follows. First, RTN amplitudes are extracted using a state-of-the-art Machine Learning (ML) algorithm together with a Hidden Markov Model (HMM) algorithm. Afterwards, their distributions across temperatures and overdrives are plotted and extensively discussed. Threshold voltage shifts due to charged defects are exponentially distributed according to different modalities at different temperatures. By fitting these distributions, the average threshold voltage shift together with the statistical defect locations are extracted. Original time traces are then used to reconstruct the low frequency noise spectra of the entire array, which follows a 1/f behaviour down to cryogenic temperatures. The LFN is higher at K, in excellent agreement with the observations already reported in literature [10, 11, 13]. The contribution to 1/f noise of defects coming from different modes and, therefore, different gate stack oxide layers, is further explored. Finally, the strong correlation between time constants and RTN amplitudes predicted by the elastic tunnelling model is tested [20].
@@ 2.1 Chip design and characterization
An array of 2500 nominally identical and minimally-spaced nanoscale nMOS devices (WxL = 90 nm x 28 nm) was used in this work (Fig.˜1(a)). It consists of 250 individually-addressable gate lines, each connected to 10 devices. The structure was designed and fabricated using a commercial 28 nm high-k metal gate CMOS technology (see Methods). The chip was bonded on a custom designed PCB and placed in a Lakeshore CPX-LVT cryogenic probe station, equipped with a custom designed 48-lines break-out box [21]. The measurements were performed on specifically designed hardware, which allowed to not only program and measure the array but also to change the sample temperature, enabling semi-automated measurement sequences.
The array was tested at six different temperatures, from T=300 K down to T=5 K. At each temperature, the characteristics in linear regime ( V) of the entire device ensemble were collected (see Fig.˜1(b)). This step was used to remove the broken devices from the analysis, extracting the threshold voltage of every working device and map the data to . The threshold voltage was calculated by using the Lambert-W function method [22], which was shown to be effective down to liquid helium temperature [23]. Afterwards, the drain current of each device was recorded for a total recording time of s, with a sampling time of ms, at five gate bias points: from to V, with a step V. The drain bias was fixed at V.
In Fig.˜1 two examples of traces showing RTN are reported. In the first case (Fig.˜1(c)), the RTN is produced by a single two-levels defect, which captures (low ) and emits (high ) individual carriers from the channel. The second case, instead, represents a more complicated signal, with two two-level defects active at the same time. A key aspect of this work is the development and the employment of an algorithm able to treat multiple defect signals. Indeed, this need turned out to be crucial for the RTN analysis at cryogenic temperatures, where the number of traces containing two defects was around 30% of the total.
To extract the defect average step height , as well as the average emission and capture time constants , a method based on a physics-informed machine learning (ML) algorithm [24] [25] coupled with a hidden Markov model (HMM) [26] was developed. The ML algorithm allowed to automatically detect RTN in the traces, distinguishing between one or multi-defect signals. In case of single-defect RTN, and were directly extracted from the signal reconstructed by the ML algorithm, whereas, in case of two active defects, the raw data were further analysed using the HMM. The benefits of this method are manifold: it requires little manual interaction, so it is suited for the treatment of large amount of data; given its complexity, it is reasonably fast (it can run on a personal computer); it allows the analysis of multi-defects RTN, as previously discussed.
@@ 2.2 Defects’ step heights distributions
In Fig.˜2(a) are plotted the CCDFs of the step heights extracted from the measurement data of approximately 2500 devices at both K and K, and V. With decreasing temperature, the shape of the CCDF changes. To determine the modality of each exponential distribution in a statistically rigorous manner, a fitting method based on the maximization of the likelihood function was developed together with an outliers detection algorithm (see Methods section). Furthermore, this fitting procedure allows the extraction of and (see eq.(1)). It is worth pointing out that includes also the undetected defects (more details can be found in the Methods section). At K the CCDF is monomodal, although the technology used in this work consists of devices with a stack. Indeed, in case of multi-layers stacks, single-defects step heights generally follow a bimodal distribution: the low- mode belongs to bulk defects, whereas the high- mode is attributed to interface defects [33, 37]. This division is based on electrostatic considerations: defects are smaller compared to ones, since they are more far away from the channel and their electrostatic impact is lower. In the present case, most of defects probably have a smaller than the detection limit of both the ML algorithm and the HMM used in the time traces analysis. At K and V, the detection limit was set to mV. It is worth pointing out that this argument finds confirmation in Fig.˜2(d), where the extracted values of are plotted. Even at K, where is generally found to be higher, as it will be discussed later, mV. Therefore, the CCDF shown in Fig.˜2(a) belongs only to defects. At this point, it can be instructive to compare the experimental with predicted by the CSA. Coherently with the literature [30, 31, 35], mV is more than two times larger than mV. As the temperature drops to K, defects start to enter the detection window, because of both a reduction of the background thermal noise and an increase of the transistors’ . Indeed, since , small threshold voltage shifts are easier to detect at low temperatures because they result in a larger . This results in the formation of a low- branch in the CCDF, which is bimodal at K (see Supplement Material).
@@
Starting from K (see Supplement Material), and down to K (Fig.˜2(a)), distributions change modality again, with the appearance of another branch at very large s. The three modes are highlighted in Fig.˜2(b), where the probability density function (PDF) of the trimodal exponential fit is plotted. This latter is given by the sum of three independent monomodal exponential PDFs, each one being described by a different set of and . The first mode is attributed to defects. The second mode is linked to defects, which enables a direct investigation of their temperature scaling from K up to K, as discussed later. The nature and origin of the third, large mode will be deeply investigated later in this section.
@@
A defect with a given amplitude is associated with each mode with a certain probability. In Fig.˜2(b), there are two values that have the same probability to belong to two modes: these are given by the intersection points and . A threshold voltage shift due to charged defects lying on the left side of , for example, is more likely to fall in the branch. The probability that the amplitude of a detected oxide defect falls into intervals of values defined by and is calculated using the cumulative distribution function (CDF) and it is plotted in Fig.˜2(c). More than 80% of measured defects have mV, and more than 60% of these belongs to . Therefore, contrary to what has been previously published on RTN at cryogenic temperatures [13] [17], most of the activity seems to come from bulk defects, which do not freeze out. In Fig.˜2(d), the temperature dependence of at V is plotted. Globally, increases at cryogenic temperatures, independently on the branch considered. Defects linked to the third mode, however, appear to be more sensitive to temperature scaling. and defects also show an increase in their average threshold voltage shift as the temperature is lowered. As it will be discussed below, this trend of can be effectively interpreted in the framework of percolation theory.
@@
In order to interpret the shape of the CCDF at cryogenic temperatures (see Fig.˜2(a) and Fig.˜2(b)), it is worth mentioning that, in the field of cryogenic electronics, percolation theory was proposed in the late ’70s to interpret the smooth change from metallic-like to thermally-activated transport in MOSFET’s inversion layer when it operates around and below and at K [38, 39, 40]. This metal-insulator transition [41] is generally attributed to surface potential fluctuations caused by interface disorder. Macroscopically, these fluctuations are usually modelled as an additional density of states below the conduction (valence) band edge [5, 7, 16]. Therefore, percolation theory has been employed independently to describe two apparently different phenomena, both originating from surface potential fluctuations: electrical conduction in band tail states at cryogenic temperatures and statistical distribution of threshold voltage shifts due to charged defects at room temperature. Starting from these considerations, percolation theory is here proposed to interpret the distributions at K (see Fig.˜2(a) and Supplement Materials). In this framework, the observed CCDFs are a manifestation of both the increased sensitivity to disorder and the appearance of interface localized states. Fluctuations in surface potential mirrors in a non uniform distribution of inversion carriers along the channel. If they have enough thermal energy, the main consequence of such fluctuations is a non homogeneous current flow from source to drain which, as previously mentioned, is responsible for the exponential distribution of s. As temperature drops, carriers thermal energy decreases accordingly, and potential fluctuations can give rise to regions where carriers are energetically prevented to flow. It is crucial to highlight that, at room temperature, the channel is modulated by disorder but it is continuous. At cryogenic temperatures, instead, the channel is a mixture of conductive paths and insulating regions [38]. In Fig.˜3(a), it is plotted a qualitative TCAD simulation performed at K to illustrate this phenomenon: at low the channel is not fully formed and conductive paths coexist together with insulating areas. The increased channel inhomogeneity amplifies the electrostatic impact of all defects at low temperatures, in agreement with the trend previously discussed and plotted in Fig.˜2(d). RTN fluctuations producing the third branch in Fig.˜2(a) are therefore interpreted as the result of the activity of those defects whose electrostatic impact is particularly enhanced by the larger channel inhomogeneity. On the other hand, especially around K, potential fluctuations can give rise to localized states, which contribute to transport through thermally-activated and also tunnelling processes. Some of the large in Fig.˜2(a) could be caused by the activity of these localized states as well, but further analyses are required to confirm that.
@@
Increasing , the modality of the CCDF changes, passing from trimodal to bimodal, as shown in Fig.˜3(c). The third branch in the CCDF vanishes, since the channel gets stronger and percolation transport is mitigated, as qualitatively illustrated in Fig.˜3(b). On the other hand, the small mode is still present. A further increase in wipes out also the branch and, at high overdrives, only the defects are visible (see Supplement Material). This can be attributed to both a reduction in measurement sensitivity and to a shift in the active energy area scanned. Concerning the first argument, at higher it is harder to detect small defects, since is larger. Moreover, changing the gate bias, the Fermi level changes as well, and the energy area of the defect bands swept changes accordingly. Therefore, at V, CCDFs are attributed only to the activity of defects in the layer.
In Fig.˜3(d), is plotted as function of . Globally, at lower temperatures is higher, independently on , as already shown in Fig.˜2(d). This is consequence of both the larger percolation and the shift in the active energy area scanned by the Fermi level at cryogenic temperatures. On the other hand, is almost independent on .